OpenStudy (anonymous):

Find the length of an arc of 40° in a circle with an 8 inch radius.

hero (hero):

associated picture?

OpenStudy (amistre64):

arc length is just a portion of circumference

OpenStudy (anonymous):

thnxx=) and no pic included

OpenStudy (amistre64):

circumference = angle * r, any arc length is angle * r

OpenStudy (amistre64):

turn the angle to rads tho

hero (hero):

Yeah, amistre has it. I was thinking about something else.

OpenStudy (amistre64):

40 needs a pi, so 40*pi/180 = rads

OpenStudy (amistre64):

and go ahead an multiply in the radius as well

hero (hero):

amistre, general formulas and approaches are always welcome

OpenStudy (amistre64):

id have to think like a pigeon then; and i tend to think on the fly lol

OpenStudy (anonymous):

i dnt get it answer choices are 8/9 , 16/9 or 64/9

OpenStudy (amistre64):

\[Arc_{length}=\frac{degree*pi*radius}{180}\]

OpenStudy (amistre64):

I get 8/9 pi when I do it

OpenStudy (anonymous):

thnxx

OpenStudy (amistre64):

40*8 = 320 320 32 16 --- = -- = -- pi; i mighta inputed it wrong in me calculator; let me dbl chk 180 18 9

OpenStudy (amistre64):

1 7/9 = 16/9 got it right that time lol

OpenStudy (anonymous):

lol ok

hero (hero):

In general, when finding the arc length of a circle when given the radius and degree, you may use the following formula: \[\frac{\theta}{360°} = \frac{x}{2 \pi r}\] , x = arc length

OpenStudy (anonymous):

thnxx guys